OpenStudy (anonymous):

If tanx= 1/2 and tany=1/3 What is tan(x+y)?

Parth (parthkohli):

\[\tan(x+y) = \dfrac{\tan x + \tan y }{1 - \tan x \tan y}\]

OpenStudy (anonymous):

\[\tan(x+y)=\frac{ \tan1/2 + \tan1/3 }{ 1- \tan1/2 \tan1/3 }\]

OpenStudy (anonymous):

Is that right?

Parth (parthkohli):

You're given that \(\tan x = 1/2\), \(\tan y = 1/3\) and not \(x = 1/2\), \(y = 1/3 \). So you will say that\[\tan(x + y) = \dfrac{1/2 + 1/3}{1 - (1/2)(1/3)}\]Do you understand?

OpenStudy (anonymous):

so you remove the tan?

OpenStudy (anonymous):

Is it cancellatin?

Parth (parthkohli):

\[\tan(x + y) = \dfrac{\tan x + \tan y }{1 - \tan x \tan y}\]Sinice you are given \(\tan x = 1/2\) and \(\tan y = 1/3\), you will replace \(\tan x\) with \(1/2\) and \(\tan y \) with \(1/3\) in the formula.

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

u just replace it with the value already given in the question